OpenStudy (hatadtadatd):

An acute angle θ is in a right triangle with sin θ = five sixths. What is the value of cot θ? Answers: five divided by the square root of eleven six divided by the square root of eleven square root of eleven divided by five square root of eleven divided by six

OpenStudy (hatadtadatd):

@phi

OpenStudy (hatadtadatd):

Can you explain more to me please

OpenStudy (phi):

Do you know about pythagorean theorem?

OpenStudy (hatadtadatd):

Yes

OpenStudy (phi):

can you write down the formula for this triangle ? fill in the letters with numbers ?

OpenStudy (hatadtadatd):

a^2+b^2=c^2

OpenStudy (phi):

yes, but put in the numbers (we know 2 out of the 3 sides, right ?)

OpenStudy (hatadtadatd):

So 5^2+6^2=c^2

OpenStudy (phi):

almost. a and b are the legs. we only know one of the legs c is the hypotenuse. we know the hypotenuse, right ?

OpenStudy (hatadtadatd):

Oh so it's 5^2+b^2=6^2?

OpenStudy (phi):

yes. it does not matter which leg (a or b) we use) next, expand 5^2 and 6^2

OpenStudy (hatadtadatd):

25+b^2=36

OpenStudy (phi):

now add -25 to both sides, and then simplify

OpenStudy (hatadtadatd):

b^2=11

OpenStudy (phi):

finally, "take the square root" of both sides the left side becomes b

OpenStudy (hatadtadatd):

That part is the part I'm confused on

OpenStudy (phi):

we get \[ b= \sqrt{11} \] we usually would leave it like that, unless we have to do calculations

OpenStudy (hatadtadatd):

Ohhhh okay so then there has to be more to it though, right?

OpenStudy (phi):

square root "undoes" the square i.e. \[ \sqrt{b^2} = b\] so you have to remember that.

OpenStudy (phi):

so \[ b^2 = 11\] do the square root on both sides \[ \sqrt{b^2} =\sqrt{11} \] and then we remember square root undoes the square: \[ b= \sqrt{11}\]

OpenStudy (hatadtadatd):

But that's not an answer choice

OpenStudy (hatadtadatd):

@phi

OpenStudy (phi):

we now know *all 3 sides* of the right triangle